OpenStudy (anonymous):

Translate this and then find the number: a number times 6 added to three is negative twenty-seven

OpenStudy (mathmale):

I'd translate this into math symbolism as follows: \[n*6 + 3 = -27, or 6n + 3=-27.\] The next step is to find a suitable solution: find whatever value of n satisfies the given equation. 6n = -27 - 3 = ? n = ?